# Attenuation of the ultra-high-energy neutrino flux by dark matter scatterings

### v1

@@ 3 Attenuation of the ultra-high energy neutrino flux
Ultra-high energy neutrinos could be produced at astrophysical sources, when protons become accelerated to extremely high energies and collide with photons surrounding the source, thereby producing charged pions that eventually decay into neutrinos. Assuming that the sources are optically thin to protons, and in view to the measured flux of ultra-high energy cosmic rays, it is possible to derive a conservative upper bound on the expected ultra-high energy neutrino flux at Earth
@@
this is the renown Waxman-Bahcall bound [22]. The actual flux is likely smaller, as the bound assumes strong redshift evolution and a proton-dominated composition of cosmic rays, among other conservative assumptions [23]. The observation of a UHE neutrino flux could then be translated into limits on the DM-$\nu$ interaction cross-section as we discuss below.
@@
The left panel of Fig. 4 summarizes the landscape of current searches for an UHE astrophysical neutrino flux. The plots shows 90% C.L. upper limits from IceCube and the Pierre Auger Observatory, obtained following [24] using the publicly available all-sky effective areas quoted in Appendix A. We also show for comparison the Waxman-Bahcall bound. It is notable that current experiments are already sensitive to UHE neutrinos from astrophysical sources.
@@
Figure 4: Left panel: Current landscape of UHE astrophysical neutrino fluxes, together with the Waxman-Bahcall upper bound. Right panel: $\chi^{2}$ of each experiment as well as the joint $\chi^{2}$ from KM3NeT, IceCube and the Pierre Auger Observatory.
@@
The figure also includes the neutrino event KM3-230213A, which is compatible with the Waxman-Bahcall bound, but however is at odds with the upper limits from IceCube and the Pierre Auger Observatory. This tension is quantified in the right-plot, which shows a Poissonian $\chi^{2}$ for the number of events, assuming that the flux is proportional to $E^{-2}$ and integrated over the KM3NeT event energy limits [0.072,2.60] EeV. As apparent from the Figure, the best fit point from combining the observation of one event at KM3NeT, and the non-observation of events at IceCube and the Pierre Auger Observatory, hints towards much lower neutrino fluxes, approximately $\sim 3\sigma$ smaller than the one reported by KM3NeT (the errors on the joint fit were obtained following the Feldman & Cousins procedure [25]). The next-generation instruments will improve the IceCube sensitivity by two orders of magnitude [26], and will elucidate whether the UHE neutrino flux saturates the Waxman-Bahcall bound, as suggested by the observation of KM3-230213A.
@@
Regardless of the true nature of KM3-230213A, current instruments already provide valuable information on the possible accelerating mechanisms at the sources, as well as on the medium through which neutrinos propagate, and specifically the DM in the intergalactic medium and in the Milky Way. In particular, the Waxman-Bahcall bound will become more stringent in the presence of sizable DM-$\nu$ interactions, as shown in Fig.5, for the concrete case when $\sigma/M_{\mathrm{DM}}=5\times 10^{-22}\,\mathrm{cm^{2}/GeV}(E_{\nu}^{0}/10^{9}\,\mathrm{GeV})^{n}$, and n=2,0,2,4n=-2,0,2,4. Conversely, the detection of a UHE neutrino flux would automatically set an upper limit on this parameter, which would depend on the energy dependence of the cross-section.
@@
Figure 5: Sky-averaged neutrino-flux, attenuated by the Milky Way DM, for an unattenuated flux saturating the Waxman-Bahcall bound. We set $\sigma/M_{\mathrm{DM}}=5\times 10^{-22}\,\mathrm{cm^{2}/GeV}(E_{\nu}^{0}/10^{9}\,\mathrm{GeV})^{n}$.
@@
Assuming that the event KM3-230213A is of diffuse origin, and requiring compatibility with the upper limits on the flux from IceCube and the Pierre Auger Observatory, one can derive an upper limit on the scattering cross-section over the mass. The effect of the attenuation on the flux is illustrated in Fig.6, for the case where the cross-section is energy independent, and considering a NFW profile. The line labeled “No attenuation” corresponds to the “Joint” fit shown in Fig. 4. As the cross-section increases, the necessary injected flux to explain the detection of the event KM3-230213A increases, and eventually would violate the Waxman-Bahcall bound. Namely, the state-of-the-art mechanisms of neutrino production in astrophysical sources are incapable of producing the event KM3-230213A, since even under the most aggressive assumptions for production, the DM-$\nu$ scatterings would attenuate the flux to unobservable levels. The upper limit reads $\sigma/M_{\rm DM}\lesssim 4\times 10^{-22}\,{\rm cm}^{2}/{\rm GeV}$. It is important to note that this limit is fairly robust against astrophysical assumptions, since the flux attenuation depends exponentially on $\sigma/M_{\rm DM}$.
@@
Figure 6: Unatennuated flux for different energy-independent DM-$\nu$ interaction cross sections, taking as a benchmark the current joint best-fit flux.
We show in Fig. 7 a compilation of the limits of the scattering cross-section over DM mass as a function of the neutrino energy. At the highest energies, we present the limit derived in this work from the assumption that KM3-230213A is of diffuse origin and is compatible with the limits from IceCube and the Pierre Auger Observatiory. At lower energies, we show the limits from the blazar TXS 0506+056, excluding [27, 28, 29] or including [8, 9] the attenuation from a DM spike surrounding the supermassive black hole at its center, the limits from the Seyfert Galaxy NGC 1068 [7], the limits from from supernova neutrinos [5, 6] and finally the limits from the Lyman-alpha forest [3]. We also show for reference the scaling of the limits when the cross-section has a simple power-law dependence with the energy. Clearly, if this dependence holds over several decades of energy, the detection of UHE neutrinos poses very stringent constraints on the DM-$\nu$ interactions at lower energies.  [^2] In many models, on the other hand, one expects a non-trivial dependence of the cross-section with the energy. For example, the scattering cross-section with a fermionic target at rest mediated by a scalar particle with mass $m_{\phi}$ goes as $E_{\nu}$ when $E_{\nu}\ll m_{\phi}$, then reaches a plateau when the center of mass energy of the collision is comparable to $m_{\phi}$, and finally decreases with the energy due to unitarity. Further, the cross-section may display resonances, akin to the Glashow resonance when a very energetic electron antineutrino collides with an electron at rest reaching a center of mass energy comparable to the $W$ mass [30].
@@
Figure 7: Limits on DM-$\nu$ scattering from the KM3-230213A event and from previous works. We assume that KM3-230213A has a diffuse origin, although the results are similar for a point source. We also show the extrapolaton of the limits to other energies assuming simple energy dependences of the cross-section.
Further, future experiments should detect many more UHE neutrinos, which will also allow the determination of their energy spectrum. The impact of the flux attenuation due to DM scatterings is shown in Fig. 8. The unattenuated neutrino flux was obtained using the public code CRPropa [31], for the 2SC-uhecr model presented in [32], and which was used in [33] as a benchmark to explain the origin of KM3-230213A. It assumes the injection of cosmic ray protons by sources with luminosity $0.12\times 10^{44}\,{\rm erg}/{\rm Mpc}^{3}/{\rm year}$ and a spectrum following a power law proportional to $E^{-0.25}$ in the energy range 5-10 EeV, and with source distribution as a function of redshift proportional to $(1+z)^{5}$ with $z_{\rm max}=4$ (left plot) or 6 (right plot), The impact of the attenuation is apparent from the plot: the peak of the distribution is shifted to lower energies, and the cut-off energy also decreases.
@@
Figure 8: Diffuse neutrino flux for different energy-dependent cross sections over DM mass: $\sigma/M_{\mathrm{DM}}=10^{-23}\mathrm{cm^{2}/GeV}(E_{\nu}^{0}/10^{8}\,\mathrm{GeV})^{n}$, for a maximum redshift of cosmic ray sources $z_{\rm max}=4$ (left plot) or (right plot). See text for details.
@@ 4 Complementarity of UHE neutrino detectors
In Section 2, we discussed the anisotropy in the neutrino flux at the surface of the Earth due to the attenuation by DM-$\nu$ scatterings in the Milky Way halo. The UHE neutrinos must then traverse the Earth before reaching the detector, which induces an additional source of anisotropy due to the attenuation of the neutrino flux by interactions with the nuclei in the interior of the Earth. The combination of different experiments at different locations. will then offer the possibility of obtaining additional information about the source of the UHE neutrino events, and possibly on the attenuation of the flux due to DM scatterings.
With current and several future detectors, the most promising candidates for ultra-high-energy neutrino events are those entering the atmosphere near the horizon. This geometry maximizes the amount of matter traversed by neutrinos, increasing the probability for them to produce a detectable particle while avoiding Earth attenuation at angles well below the horizon. Therefore, experiments “sweeping” the sky as the horizon moves against the fixed stars are well-suited for the detection of UHE neutrino flux generated by distant sources.
@@
We show in Fig. 9 the day averaged effective areas (i.e integrating over the right ascension) of IceCube, the Pierre Auger Observatory and KM3NeT as a function of the declination. KM3NeT (located at a latitude of $36.27^{\circ}$ N) has therefore sensitivity to a wide range of declinations, being only blind to UHE neutrinos arriving with $90^{\circ}\leq\delta\lesssim 36^{\circ}$, since this region of the sky remains below the horizon over the whole year. IceCube (located at a latitude of $90^{\circ}$ S) is practically blind to neutrinos arriving from positive declinations, and presents a high peak of sensitivity at $\delta\simeq 0$, since this direction lies at the horizon of the South Pole the whole year. Finally, the Pierre Auger Observatory (located at a latitude of $35.0^{\circ}$ S) is only sensitive to neutrinos from a very narrow range of incoming directions. Although the effect is partly alleviated by the rotation of the Earth, the day-averaged effective area is still quite angle-dependent. Notice that the effective area of KM3NeT is a factor of $\sim 10$ smaller than the one from IceCube, even though they are both km3 experiments, the reason being that KM3NeT is still under construction.
@@
Figure 9: Day-averaged effective area of IceCube, the Pierre Auger Observatory and KM3NeT, as a function of the declination of arrival of the neutrino flux. We also show as comparison the expected dependence of the declination of a diffuse UHE neutrino flux for different values of the scattering cross-section over mass, due to the attenuation of the neutrino flux in the Milky Way halo assuming a NFW profile.
The figure also shows the expected astrophysical UHE neutrino flux for different values of the scattering cross-section over DM mass. Clearly, for an unattenuated flux (purely isotropic) IceCube has a very good sensitivity. On the other hand, if the flux is attenuated by DM scatterings in the Milky Way halo, the flux at negative declinations decreases, and IceCube loses sensitivity. Other instruments, such as KM3NeT or Auger, which are sensitive to UHE neutrinos coming from positive declinations, may observe signals while escaping detection at IceCube.
This is further illustrated in Fig. 10, which shows the ratios of time-integrated effective areas of IceCube and KM3NeT (green line), and Auger and KM3NeT (orange line) as a function of $\sigma/M_{\rm DM}$. The relative exposures of Auger and KM3NeT are roughly independent of the value of $\sigma/M_{\rm DM}$, due to their locations on the Earth. On the other hand, for small cross-sections IceCube has a much bigger effective relative to KM3NeT, and as the cross-section increases, the effective area relative to KM3NeT decreases. This non-trivial interplay of the effective areas of both experiments could in principle open the intriguing possibility that KM3Net could observe UHE neutrino events, while escaping detection at IceCube, and thereby potentially explaining the observation of KM3-230213A. Quantitatively, however, the difference in effective areas without attenuation is so large that the impact of DM attenuation is mild: using a saturated Poisson likelihood test [34], we find that the tension between all three experiments can get reduced from $\sim 3\sigma$ [11] to, at most, $\sim 2.8\sigma$. Nevertheless, this is an important effect to keep in mind should further observations of UHE neutrinos with comparable detectors be in mild apparent tension among them.
@@
Figure 10: Effective sensitivity of IceCube and the Pierre Auger Observatory, relative to KM3NeT, of a diffuse UHE neutrino flux as a function of the scattering cross-section over DM mass, due to the interplay of the effective area and arrival direction of the neutrino flux attenuated by DM scatterings in the MW halo assuming a NFW profile.

### v2

@@ 3 Attenuation of the ultra-high energy neutrino flux
In principle, using the observed UHE neutrino flux to explore the neutrino-DM cross section requires a priori assumptions on the flux before attenuation. Here, we discuss how mild astrophysical assumptions allow to set conservative bounds.
Ultra-high energy neutrinos could be produced at astrophysical sources, when protons get accelerated to extremely high energies and collide with photons surrounding the source, thereby producing charged pions that eventually decay into neutrinos. Assuming that the sources are optically thin to protons, and in view to the measured flux of ultra-high energy cosmic rays, it is possible to derive a conservative upper bound on the expected ultra-high energy neutrino flux at Earth
@@
this is the renowned Waxman-Bahcall bound [26]. The actual flux is likely smaller, as the bound assumes strong redshift evolution and a proton-dominated composition of cosmic rays, among other conservative assumptions [27]. The observation of a UHE neutrino flux could then be translated into limits on the DM-$\nu$ interaction cross section as we discuss below.
@@
The left panel of Fig. 4 summarizes the landscape of current searches for an UHE astrophysical neutrino flux. The plots shows the latest 90% C.L. upper limits from IceCube and the Pierre Auger Observatory, obtained following Ref. [28] using the publicly available all-sky effective areas from Refs. [13, 12, 11] as described in Appendix A. We also show for comparison the Waxman-Bahcall bound. It is notable that current experiments are already sensitive to UHE neutrinos from astrophysical sources.
@@
Figure 4: Left panel: Current landscape of UHE astrophysical neutrino fluxes, together with the Waxman-Bahcall upper bound. All limits and measurements are at 90% CL. Right panel: $\chi^{2}$ of each experiment as well as the joint $\chi^{2}$ from KM3NeT, IceCube and the Pierre Auger Observatory.
@@
The figure also includes the neutrino event KM3-230213A, which is compatible with the Waxman-Bahcall bound, but it is however at odds with the upper limits from IceCube and the Pierre Auger Observatory. This tension is quantified in the right-plot, which shows a Poissonian $\chi^{2}$ for the number of events, assuming that the flux is proportional to $E^{-2}$ and integrating over the 90% CL energy limits of the KM3NeT event, [0.072,2.60] EeV. As apparent from the Figure, the best-fit point from combining the observation of one event at KM3NeT and the non-observation of events at IceCube and the Pierre Auger Observatory hints towards much lower neutrino fluxes, approximately $\sim 3\sigma$ smaller than the one reported by KM3NeT [29, 30, 12]. We obtain the errors on the joint fit following the Feldman & Cousins procedure [31], analogously to Ref. [30] (we have checked that if we use the same IceCube dataset as Ref. [30] we reproduce their results). Next-generation instruments will improve the IceCube sensitivity by two orders of magnitude [32], and will elucidate whether the UHE neutrino flux saturates the Waxman-Bahcall bound, as suggested by the observation of KM3-230213A.
@@
Regardless of the true nature of KM3-230213A, current instruments already provide valuable information on the possible accelerating mechanisms at the sources, as well as on the medium through which neutrinos propagate, and specifically the DM in the intergalactic medium and in the Milky Way. In particular, if the unattenuated flux satisfies the Waxman-Bahcall bound, in the presence of sizable DM-$\nu$ interactions the bound on the attenuated flux at Earth will be more stringent. We show this in Fig. 5 for the specific case of $\sigma/M_{\mathrm{DM}}=5\times 10^{-22}\,\mathrm{cm^{2}/GeV}\times(E_{\nu}/10^{9}\,\mathrm{GeV})^{n}$ and n=2,0,2,4n=-2,0,2,4; where we have conservatively included only attenuation due to the Milky Way. Conversely, the detection of a UHE neutrino flux would automatically set an upper limit on $\sigma/M_{\mathrm{DM}}$, which would depend on the energy dependence of the cross section.
@@
Figure 5: Sky-averaged neutrino-flux, attenuated by the Milky Way DM, for an unattenuated flux saturating the Waxman-Bahcall bound. We set $\sigma/M_{\mathrm{DM}}=5\times 10^{-22}\,\mathrm{cm^{2}/GeV}\times\left(\frac{E_{\nu}}{10^{9}\,\mathrm{GeV}}\right)^{n}$.
@@
Assuming that the event KM3-230213A is of diffuse origin, and inferring the corresponding flux by combining with the upper limits from IceCube and the Pierre Auger Observatory, one can derive an upper limit on the scattering cross section over the mass. The effect of the attenuation on the flux is illustrated in Fig. 6, for an energy-independent cross section and considering an NFW profile. The line labeled “No attenuation” corresponds to the “Joint” fit shown in Fig. 4. As the cross section increases, the necessary injected flux to explain the detection of the event KM3-230213A increases, and it would eventually violate the Waxman-Bahcall bound. Namely, the state-of-the-art mechanisms of neutrino production in astrophysical sources are incapable of producing the event KM3-230213A, since even under the most aggressive assumptions for production, the DM-$\nu$ scatterings would attenuate the flux to unobservable levels. The upper limit reads $\sigma/M_{\rm DM}\lesssim 4\times 10^{-22}\,{\rm cm}^{2}/{\rm GeV}$. It is important to note that this limit is fairly robust against astrophysical assumptions, since the flux attenuation depends exponentially on $\sigma/M_{\rm DM}$. This is also illustrated in Fig. 8: if $\sigma/M_{\mathrm{DM}}$ is larger than our limit by a factor $\sim 2$, the unattenuated flux should violate the Waxman-Bahcall bound by about two orders of magnitude. On top of that, we conservatively include only attenuation by the Milky Way.
@@
Figure 6: Unatennuated flux for different energy-independent DM-neutrino interaction cross sections, taking as a benchmark the current joint best-fit flux.
@@
Figure 7: Limits on DM-$\nu$ scattering from the KM3-230213A event and from previous works. We assume that KM3-230213A has a diffuse origin, although the results are similar for a point source (under different assumptions, see text). We also show the extrapolation of the limits to other energies assuming simple energy dependence of the cross section.
We show in Fig. 7 a compilation of the limits on the scattering cross section over DM mass as a function of the neutrino energy. At the highest energies, we present the limit derived in this work from the assumption that KM3-230213A is of diffuse origin (similar limits can be obtained if the event originates from a point source and exponentially large attenuation is a priori excluded). At lower energies, we show the limits from the blazar TXS 0506+056, excluding [33, 34, 35] or including [9, 10] the attenuation from a DM spike surrounding the supermassive black hole at its center, the limits from the Seyfert Galaxy NGC 1068 [8], the limits from from supernova neutrinos [6, 7] and finally the limits from Milky Way satellite galaxies [5]. We also show for reference the scaling of the limits when the cross section has a simple power-law dependence with neutrino energy. Clearly, if this dependence holds over several decades of energy, the detection of UHE neutrinos poses very stringent constraints on DM-neutrino interactions at lower energies.
In many models, one expects indeed a non-trivial dependence of the cross section with energy (see Ref. [25] for explicit examples, and Ref. [36] for a critical overview of models with observable cross sections). For example, the scattering cross section with a fermionic target at rest mediated by a scalar particle with mass $m_{\phi}$ grows as $E_{\nu}$ when $E_{\nu}\ll m_{\phi}$ if the target mass is much smaller than $E_{\nu}$, or as $E_{\nu}^{2}$ if the target mass is much higher than $E_{\nu}$. It then reaches a plateau when the center-of-mass energy of the collision is comparable to $m_{\phi}$, and it finally decreases with energy due to unitarity. As another example, the cross section with a scalar target mediated by a heavy fermion grows as $E_{\nu}^{3}$ if the target mass is much smaller than $E_{\nu}$, or as $E_{\nu}^{4}$ if it is much larger. Further, the cross section may display resonances, akin to the Glashow resonance when a very energetic electron antineutrino collides with an electron at rest reaching a center-of-mass energy comparable to the $W$ mass [37]. Thus, depending on the particular model, UHE neutrino attenuation may be the leading probe of DM-$\nu$ interactions.
@@
Figure 8: Diffuse neutrino flux for different energy-dependent cross sections over DM mass: $\sigma/M_{\mathrm{DM}}=10^{-23}\mathrm{cm^{2}/GeV}\times(E_{\nu}^{0}/10^{8}\,\mathrm{GeV})^{n}$, assuming a particular model of UHE neutrino production and for a maximum redshift of cosmic ray sources $z_{\rm max}=4$ (left plot) or (right plot). See text for details.
Further, future experiments should detect many more UHE neutrinos, which will also allow to determine their energy spectrum. The impact of the flux attenuation due to DM scatterings is shown in Fig. 8. We illustrate the effect with an unattenuated neutrino flux obtained using the public code CRPropa [38], for the 2SC-uhecr model presented in [39], and which was used in Ref. [40] as a benchmark to explain the origin of KM3-230213A. It assumes the injection of cosmic-ray protons by sources with luminosity $0.12\times 10^{44}\,{\rm erg}/{\rm Mpc}^{3}/{\rm year}$, a spectrum following a power law proportional to $E^{-0.25}$ in the energy range 5-10 EeV, and a source distribution as a function of redshift proportional to $(1+z)^{5}$ up to a maximum redshift $z_{\rm max}=4$ (left plot) or 6 (right plot). The impact of the attenuation and its energy dependence is apparent from the plot: the peak of the distribution is shifted to lower energies, and the cut-off energy also decreases. The energy dependence of the cross section is also relevant, as an enhanced cross section at high energies suppresses the flux of neutrinos produced at high redshift, which for the same present-day energy had higher energies at production.
This plot illustrates that very aggressive astrophysical assumptions in the form of an assumed cosmic-ray spectrum, chemical composition, and redshift-dependent injection; would allow to tightly probe neutrino-DM interactions. More importantly, it highlights the relevance of constraining neutrino-DM interactions, as one of the main astrophysics goals of UHE neutrino detection is to learn about the sources, chemical composition, and redshift evolution of UHE cosmic rays. As the figure shows, this goal can be hindered by neutrino-DM interactions.
@@ 4 Complementarity of UHE neutrino detectors
In Section 2 above, we have discussed the anisotropy in the neutrino flux at the surface of the Earth due to the attenuation by scattering with DM in the Milky Way halo. Testing this anisotropy would be a very robust probe of neutrino-DM scattering. This feature has been exploited with high-energy astrophysical neutrino data [24, 25] but, as we discuss below, requires a different approach at UHE energies.
Before reaching the detector, UHE neutrinos must traverse the Earth. This attenuates the neutrino flux by interactions with the nuclei at Earth. Thus, different experiments are generically sensitive to different fractions of the sky. The combination of different experiments at different locations is hence mandatory to look for anisotropies induced by DM-$\nu$ scattering.
@@
With current and several future detectors, the UHE neutrino sensitivity is maximized near the horizon. This enlarges the amount of matter traversed by neutrinos, increasing the probability for them to produce a detectable interaction, while avoiding Earth attenuation at angles well below the horizon. Therefore, experiments “sweeping” the sky as the horizon moves against the fixed stars are well-suited to be sensitive at different directions in the sky. We show in Fig. 9 the day-averaged effective areas (i.e., integrating over the right ascension) of IceCube, the Pierre Auger Observatory and KM3NeT as a function of the declination. KM3NeT (located at a latitude of $36.27^{\circ}$ N) has sensitivity to a wide range of declinations, being only blind to UHE neutrinos arriving with $90^{\circ}\leq\delta\lesssim 36^{\circ}$, since this region of the sky remains below the horizon over the whole day. IceCube(located at a latitude of $90^{\circ}$ S) is, in contrast, practically blind to neutrinos arriving from positive declinations, and presents a high peak of sensitivity at $\delta\simeq 0$, since this direction lies at the horizon of the South Pole the whole day. Finally, the Pierre Auger Observatory (located at a latitude of $35.0^{\circ}$ S) is only sensitive to neutrinos from a very narrow range of incoming directions [13]. Although the effect is partly alleviated by the rotation of the Earth, the day-averaged effective area is still rather angle-dependent. Notice that the effective area of KM3NeT is a factor of $\sim 10$ smaller than the one from IceCube, even though they are both km3 experiments, the reason being that KM3NeT is still under construction.
@@
Figure 9: Day-averaged effective area of IceCube, the Pierre Auger Observatory and KM3NeT, as a function of the declination of the arriving neutrino. We also show as comparison the expected angular dependence of a diffuse UHE neutrino flux for different scattering cross sections over DM mass, due to attenuation in the Milky Way halo and assuming an NFW profile.
@@
Figure 10: Effective sensitivity of IceCube and the Pierre Auger Observatory, relative to KM3NeT, of a diffuse UHE neutrino flux as a function of the scattering cross section over DM mass, due to the interplay of the effective area and arrival direction of the neutrino flux attenuated by DM scatterings in the MW halo assuming a NFW profile.
The figure also shows the expected astrophysical UHE neutrino flux for different values of the scattering cross section over DM mass. Clearly, for an unattenuated flux (purely isotropic) IceCube has a very good sensitivity. In turn, if the flux is attenuated by scattering with DM in the Milky Way halo, the flux at negative declinations decreases (see Fig. 2 for the angular dependence of the DM column density), and IceCube loses sensitivity. Other instruments, such as KM3NeT or Auger, that are sensitive to UHE neutrinos coming from positive declinations, may observe signals while escaping detection at IceCube.
This is further illustrated in Fig. 10, which shows the ratios among the effective areas of IceCube and KM3NeT (green line), and Auger and KM3NeT (orange line) as a function of $\sigma/M_{\rm DM}$ and integrated over the different exposure times of the detectors. The relative exposures of Auger and KM3NeT are roughly independent of the value of $\sigma/M_{\rm DM}$, due to their locations on Earth. In turn, for small cross sections IceCube has a much bigger effective area relative to KM3NeT, and as the cross section increases, the effective area relative to KM3NeT decreases because IceCube is most sensitive to the region in the sky with the largest attenuation. This non-trivial interplay of the effective areas of both experiments could in principle open the intriguing possibility that KM3Net could observe UHE neutrino events, while escaping detection at IceCube, and thereby potentially explain the observation of KM3‑230213A. Quantitatively, however, the difference in effective areas without attenuation is so large that the impact of DM attenuation is mild: using a saturated Poisson likelihood test [41], we find that the tension between all three experiments can reduce from $\sim 3\sigma$ [29, 30, 12] to, at most, $\sim 2.8\sigma$. Nevertheless, this is an important effect to keep in mind should further observations of UHE neutrinos with comparable detectors be in mild apparent tension among them.